how would i make a system of linear equations for the following problem when i want to use either the elimination or substitution method?
On a canoe trip, Rita paddled upstream (against the current) at an average speed of 2 mi/h relative to the river bank. on the return trip downstream (with the current), her average speed was 3 mi/h. find rita's paddling speed in still water and the speed of the river's current.

Answer by ankor@dixie-net.com(22740)

On a canoe trip, Rita paddled upstream (against the current) at an average speed of 2 mi/h relative to the river bank. on the return trip downstream (with the current), her average speed was 3 mi/h. find rita's paddling speed in still water and the speed of the river's current.
:
Let x = paddling speed in still water; y = speed of the current;
:
Make two equations, finding the average speed as given:
(x - y)/2 = 2
(x + y)/2 = 3
Get rid of the denominators mult both equations by 2, use the elimination method
x - y = 4
x + y = 6
---------- add
2x + 0 = 10
x = 10/2
x = 5 mph
:
Find y: 5 - y = 4; y = 1 mph
:
Check: (5+1)/2 = 3 mph and (5-1)/2 = 2